Gucchi:

Simplify the square root of four over the cubed root of four. four raised to the five sixths power four raised to the one sixth power four raised to the three halves power four raised to the one third power

Gucchi:

am i right with B?

snowflake0531:

\[\frac{ \sqrt{4}}{\sqrt[3]{4} }\] \[\sqrt{4} = 4^{\frac{1}{2}}\] \[\sqrt[3]{4}=4^{\frac{1}{3}}\] \[\frac{a^x}{a^y}=a^{x-y}\] \[\frac{ \sqrt{4}}{\sqrt[3]{4} } = \frac{ 4^{\frac{1}{2}}}{ 4^{\frac{1}{3}} }=4^{\frac{1}{2}-\frac{1}{3}}=4^{\frac{3}{6}-\frac{2}{6}}=4^\frac{1}{6}\] so yeah, you're correct

Gucchi:

THANK YOU!

snowflake0531:

yw~